What maximum height attained by the object that is projected vertically upward with a
speed of 60 m/s? How long will it take for the object to reach the maximum height?
1
Expert's answer
2020-10-19T13:24:48-0400

Let's first find the time the object takes to reach the maximum height from the kinematic equation:


"v=v_0-gt,""0=v_0-gt,""t=\dfrac{v_0}{g}."

Then, we can substitute "t" into another kinematic equation and find "y_{max}":


"y_{max}=v_0t-\dfrac{1}{2}gt^2,""y_{max}=v_0\dfrac{v_0}{g}-\dfrac{1}{2}g(\dfrac{v_0}{g})^2,""y_{max}=\dfrac{v_0^2}{2g}=\dfrac{(60\ \dfrac{m}{s})^2}{2\cdot 9.8\ \dfrac{m}{s^2}}=183.7\ m."

Finally, we can find the time the object takes to reach the maximum height:


"t=\dfrac{v_0}{g}=\dfrac{60\ \dfrac{m}{s}}{9.8\ \dfrac{m}{s^2}}=6.1\ s."

Answer:

"y_{max}=183.7\ m, t=6.1\ s."